OpenStudy (anonymous):

What is the next term in the number pattern? 100, 75, 50, 25 is the next number 5?

OpenStudy (anonymous):

i wouldn't think so, looks like you are subtracting 25 each time

OpenStudy (anonymous):

because its going by 15

OpenStudy (anonymous):

i don't see a 15 in it

OpenStudy (anonymous):

oh my options are 0 5 -25 1

OpenStudy (anonymous):

\(100-25=75\) \(75-25=50\) \(50-25=25\)

OpenStudy (anonymous):

last one would be \(25-25\)

OpenStudy (anonymous):

0